A particle starts at the origin and moves along the positive x-axes such that its distance from the origin at any time is determined by the equation: x=2t^3 -9t^2 +12t Find the 3.1) times when the velocity is zero. 3.2) acceleration when the velocity is zero.


1
Expert's answer
2021-07-05T09:12:42-0400

3.1) times when the velocity is zero.

we are given the equation "X=2t^3 -9t^2 +12t"

we need to differentiate this function with respect to t to obtain velocity equation

"\implies" "{dX\over dt}=V(t)=6t^2-18t+12"

we are asked to find times when velocity is zero

"\implies" "V(t)=6t^2-18t+12=0"

divide by 6 through

"{6t^2\over 6}-{18t\over 6}+{12\over 6}=0"

"\implies" "(t-1)(t-2)=0"

"\therefore t=1" sec and "t=2" secs

3.2) acceleration when the velocity is zero.

To get acceleration, we differentiate velocity equation with respect to time

"\implies{dV\over dt}=a(t)=12t-18"

when "t=2"

"\implies a(2)=12(2)-18=6" "m\/s^2"

When "t=1"

"\implies a(1)=12(1)-18=-6" "m\/s^2"

"\therefore" decelarition when t=1 is 6 "m\/s^2"